### v1

# bibliography.bib

Abstract
We draw from Raney duality and generalize the notion of canonical extension for distributive lattices to the context of frames. This is a generalization because for distributive lattices we are in the context of coherent spaces, whereas frames represent arbitrary spaces. We introduce Raney extensions of frames, pairs $(L,C)$ where $C$ is a coframe, $L\subseteq C$ is a frame that meet-generates it, and the inclusion $L\subseteq C$ preserves the frame operations as well as the strongly exact meets. We think of these as being algebraic versions of the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ of a frame of opens into a lattice of saturated sets (upper sets in the specialization order). We show that a frame may have several Raney extensions, and all satisfy a generalization of the properties of density and compactness from the theory of canonical extensions. We show that every frame $L$ has the largest and the smallest Raney extension, and that these are, respectively, the coframe of fitted sublocales $\mathsf{S}_{\mathfrak{o}}(L)$ and the opposite of the frame $\mathsf{S}_{\mathfrak{c}}(L)$ of joins of closed sublocales (both studied in [moshier20] and [moshier22]). We thus show that these structures have universal properties which are dual of one another. For Raney extensions $(L,C)$ and $(M,D)$, we characterize the frame morphisms $f:L\to M$ which can be extended to a morphism of Raney extensions. We apply this result to obtain a characterization of morphisms of frames $f:L\to M$ which can be lifted to frame morphisms $\mathsf{S}_{\mathfrak{c}}(L)\to\mathsf{S}_{\mathfrak{c}}(M)$. We also show that the canonical extension of a locally compact frame, as defined in [jakl20], is the free Raney extension over it such that it is algebraic (generated by its compact elements). We give a characterization of sobriety and of strict sobriety based on a variation of the compactness property of the extension $(\Omega(X),\mathcal{U}(X))$.
1 Introduction
In this work, we study an algebraic, pointfree version of the embedding of a frame of open sets into a lattice of saturated sets. Our constructions are inspired by Raney duality, as illustrated in [bezhanishvili20]. In Raney duality, rather than mapping a space $X$ to the frame $\Omega(X)$ of its open sets as one does in pointfree topology, we map it to the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ of its open sets into the lattice of saturated sets [^1] Saturated sets are intersections of open sets.. In this duality, on the algebraic side all objects of the category have, so to speak, enough points. That is, every Raney algebra is of the form $(\Omega(X),\mathcal{U}(X))$ for some space $X$. In this work, we extend the category of Raney algebras to include pointfree objects. We also build on the notion of canonical extension of a distributive lattice: an algebraic, pointfree version of the embedding of the lattice of compact open sets of a coherent space into the lattice of saturated sets. For a distributive lattice, its canonical extension is unique. For frames, we claim, there is no unique way of extending a frame $L$ to a pointfree lattice of saturated sets.
We will consider as objects of our category Raney extensions, pairs $(L,C)$ where $C$ is a coframe and $L\subseteq C$ is a frame which meet-generates $C$ and such that the embedding preserves the frame operations together with strongly exact meets [^2] Strongly exact meets are the pointfree version of those intersections of open sets which are open. Because a meet of a collection $\{U_{i}:i\in I\}$ of opens in general is calculated as the interior of $\bigcap_{i}U_{i}$, these are exactly the meets that are preserved by the embedding $\Omega(X)\subseteq\mathcal{U}(X)$.. We use classical results from the theory of polarities of Birkhoff to show that all Raney extensions are, up to isomorphism, pairs of the form $(L,\mathcal{F}^{op})$, where $\mathcal{F}\subseteq\mathsf{Filt}(L)$ is a sublocale of the collection of filters of $L$. We build on results in [jakl24] to show that all Raney extensions have universal properties which are variations of the universal property of the canonical extension, by considering variations of the density and compactness property. In particular, for a collection of filters $\mathcal{F}\subseteq L$ of a frame $L$ we will say that a Raney extension $(L,C)$ is
  • dense with respect to $\mathcal{F}$ if every element of $C$ is a join of elements of the form $\bigwedge F$ for $F\in\mathcal{F}$;
  • compact with respect to $\mathcal{F}$ if for each $F\in\mathcal{F}$ and each $a\in L$ we have that $\bigwedge F\leq a$ implies $a\in F$.
We also tackle the problem of lifting of morphisms. For Raney extensions $(L,C)$ and $(M,D)$, we will characterize frame morphisms $f:L\to M$ such that can be extended to morphisms of Raney extensions. Recently, the collection $\mathsf{S}_{\mathfrak{c}}(L)$ has received quite a lot of attention in pointfree topology. See [ball18], [joinsofclosed19], [moshier20], [ball19], [ball16]. This is the collection of joins of closed sublocales. Functoriality of the assignment $L\mapsto\mathsf{S}_{\mathfrak{c}}(L)$ is explored in [ball19] for subfit frames, and the authors identify some ways of restricting the objects so that all frame morphisms $f:L\rightarrow M$ lift to frame morphisms. We will use our characterization on how frame maps can be extended to maps of Raney extensions to find conditions on morphisms $f:L\rightarrow M$ between arbitrary frames (not necessarily subfit) to lift, and find that a morphism lifts if and only if for an exact meet $\bigwedge_{i}x_{i}\in L$ we have both that $\bigwedge_{i}f(x_{i})$ is exact and that it equals $f(\bigwedge_{i}x_{i})$.
Another important structure in pointfree topology is the subcoframe $\mathsf{S}_{\mathfrak{o}}(L)\subseteq\mathsf{S}(L)$ of fitted sublocales, that is, intersections of open sublocales. See, for instance, [clementino18], [moshier20]. The structures $\mathsf{S}_{\mathfrak{c}}(L)$ and $\mathsf{S}_{\mathfrak{o}}(L)$ are compared in [moshier20] and in [moshier22]. In this work we show that, for a frame $L$, $\mathsf{S}_{\mathfrak{o}}(L)$ and $\mathsf{S}_{\mathfrak{c}}(L)$ are, respectively, the largest and smallest Raney extensions, showing that these two structures have universal properties which are dual to each other. For $\mathsf{S}_{\mathfrak{o}}(L)$ it is also true that it is the free Raney extension over $L$.
Finally, we will look at canonical extensions of frames, as defined in [jakl20], and show that for a pre-spatial frame its canonical extension is the free Raney extension over it which is algebraic (join-generated by its compact elements), thus showing another universal property of this construction. We will also show that a $T_{0}$ space is sober, resp. strictly sober, if and only if the Raney extension $(\Omega(X),\mathcal{U}(X))$ is compact with respect to the collection of completely prime, resp. Scott-open, filters.
2 Background
2.1 Sublocales
We will work in the category $\mathbf{Frm}$. Sometimes in pointfree topology one works in the category $\mathbf{Loc}$ of locales. The category $\mathbf{Loc}$ is defined as the category whose objects are frames, which are referred to as locales when adopting this approach. The morphisms of $\mathbf{Loc}$ are the right adjoints to frame maps. Hence, a frame map $f:L\rightarrow M$ will correspond to the morphism $f_{*}:M\rightarrow L$ in $\mathbf{Loc}$. The category $\mathbf{Loc}$ is dually isomorphic to $\mathbf{Frm}$. In the category of topological spaces, subspace inclusions are, up to isomorphism, the regular monomorphisms. A sublocale is a regular monomorphism in $\mathbf{Loc}$. Even when working with frames, the term sublocale is still used. We follow Picado and Pultr in [picadopultr2012frames] in defining a sublocale of a frame $L$ to be a subset $S\subseteq L$ such that:
  1. 1.
    It is closed under all meets;
  2. 2.
    Whenever $s\in S$ and $x\in L$ we have $x\rightarrow s\in S$.
These requirements are equivalent to stating that $S\subseteq L$ is a regular monomorphism in $\mathbf{Loc}$. Observe that the collection of sublocales of a frame is closed under all intersections. We have the following useful fact.
Lemma 2.1.
If $S$ and $T$ are sublocales of $L$ such that $S\subseteq T$, then $S$ is a sublocale of $T$.
The family $\mathsf{S}(L)$ of all sublocales of $L$ ordered by inclusion is a coframe. Meets in $\mathsf{S}(L)$ are set-theoretical intersections. The top element is $L$ and the bottom element is $\{1\}$. Because $\mathsf{S}(L)$ is a coframe, there is a difference operator on it, dual to Heyting implication, defined for sublocales $S$ and $T$ as $S{\setminus}T=\bigcap\{U\in\mathsf{S}(L):S\subseteq T\cup U\}$. For a sublocale $S$, we denote the element $L{\setminus}S$ as $S^{*}$, and we call it the supplement of $S$.
Open and closed sublocales
We now focus on the pointfree versions of open and closed subspaces. For each $a\in L$, there are an open sublocale and a closed sublocale associated with it. These are, respectively,
$\mathfrak{o}(a)=\{a\rightarrow b:b\in L\},$ $\mathfrak{c}(a)=\uparrow a.$
Open and closed sublocales behave like open and closed subspaces in many respects; we list a few of them below.
Proposition 2.2.
For every frame $L$ and $a,b,a_{i}\in L$ we have
  1. 1.
    $\mathfrak{o}(1)=L$ and $\mathfrak{o}(0)=\{1\}$;
  2. 2.
    $\mathfrak{c}(1)=\{1\}$ and $\mathfrak{c}(0)=L$;
  3. 3.
    $\bigvee_{i}\mathfrak{o}(a_{i})=\mathfrak{o}(\bigvee_{i}a_{i})$ and $\mathfrak{o}(a)\cap\mathfrak{o}(b)=\mathfrak{o}(a\wedge b)$;
  4. 4.
    $\bigcap_{i}\mathfrak{c}(a_{i})=\mathfrak{c}(\bigwedge_{i}a_{i})$ and $\mathfrak{c}(a)\vee\mathfrak{c}(b)=\mathfrak{c}(a\wedge b)$;
  5. 5.
    The elements $\mathfrak{o}(a)$ and $\mathfrak{c}(a)$ are complements of each other in $\mathsf{S}(L)$: we have $\mathfrak{o}(a)\cap\mathfrak{c}(a)=$ and $\mathfrak{o}(a)\vee\mathfrak{c}(a)=L$;
  6. 6.
    $\mathfrak{c}(a)\subseteq\mathfrak{o}(b)$ if and only if $a\vee b=1$, and $\mathfrak{o}(a)\subseteq\mathfrak{c}(b)$ if and only if $a\wedge b=0$.
The ordered collection of closed sublocales of a frame $L$ is a subcoframe of $\mathsf{S}(L)$. In fact, this coframe is anti-isomorphic to $L$. The collection of joins of closed sublocales of a frame is a frame and is called $\mathsf{S}_{\mathfrak{c}}(L)$. Open sublocales form a subframe of $\mathsf{S}(L)$, and this is isomorphic to $L$. The sublocales which are intersections of open sublocales are called fitted. These form a subcolocale of $\mathsf{S}(L)$ which is called $\mathsf{S}_{\mathfrak{o}}(L)$.
2.2 Saturated sets and fitted sublocales
In a topological space $X$, we can define the specialization preorder on its points, defined as $x\leq y$ whenever $x\in U$ implies $y\in U$ for all open sets $U\subseteq X$. In this paper, for a space $X$, we will denote as $\mathcal{U}(X)$ the ordered collection of all upper sets in the specialization preorder, and for a point $x\in X$ we denote as ${\uparrow}x$ the upper set of $x$ with respect to this preorder. A space $X$ is $T_{0}$ if and only if the specialization preorder is an order. A space is $T_{1}$ if and only if the specialization order on it is discrete. For a space $X$, we denote as $\mathcal{U}(X)$ the lattice of its upsets (upper-closed sets) under the specialization preorder. The following is a standard fact of topology, which can be easily checked.
Proposition 2.3.
For a topological space $X$, a subset is an upset in the specialization preorder if and only if it is saturated.
The following is an important theorem by Hofmann and Mislove. A filter of a frame $L$ is Scott-open if it is not accessible by directed joins. We call $\mathsf{Filt}_{\mathcal{SO}}(L)$ the ordered collection of Scott-open filters of a frame $L$.
Theorem 2.4.
([hofmann81], Theorem 2.16) If the Prime Ideal Theorem holds, then for each sober space $X$ there is an anti-isomorphism between $\mathsf{Filt}_{\mathcal{SO}}(\Omega(X))$ and the ordered collection of compact saturated sets of $X$, assigning to each filter $F$ the set $\bigcap F$.
Exact and strongly exact filters
[…]
$F\mapsto\bigcap\{\mathfrak{o}(f):f\in F\}.$
This maximally restricts to an anti-isomorphism between fitted sublocales and strongly exact filters (see [moshier20]. Recall that a meet $\bigwedge_{i}x_{i}$ is strongly exact if for all $y\in L$ we have that $x_{i}\rightarrow y=y$ implies $(\bigwedge_{i}x_{i})\rightarrow y=y$, and that a filter is strongly exact if it is closed under strongly exact meets. We call $\mathsf{Filt}_{\mathcal{SE}}(L)$ the ordered collection of strongly exact filters. This is a frame where meets are computed as intersections, and additionally it is a sublocale of $\mathsf{Filt}(L)$.
Lemma 2.5.
For a filter $F\subseteq L$, the following are equivalent.
  1. 1.
    $F$ is strongly exact.
  2. 2.
    $\bigcap_{f\in F}\mathfrak{o}(f)\subseteq\mathfrak{o}(x)$ is equivalent to $x\in F$.
  3. 3.
[…]
  4. 4.
    $F=\{x\in L:S\subseteq\mathfrak{o}(x)\}$ for some sublocale $S\subseteq L$.
[…]
Theorem 2.6.
We have an isomorphism of coframes
[…]
$S\mapsto\{a\in A:S\subseteq\mathfrak{o}(a)\}.$
We also have an isomorphism of frames
$\mbox{coker}:\mathsf{S}_{\mathfrak{c}}(L)\cong\mathsf{Filt}_{\mathcal{E}}(L),$
$S\mapsto\{a\in L:\mathfrak{c}(a)\subseteq S\}.$
Notable collections of filters
In this paragraph, we refer to [jakl24], and mention the main results from that we are going to use. We will refer to several important concrete collections of filters. Since the collection $\mathsf{Filt}(L)$ is a frame, there is a Heyting operation $\rightarrow$ on it. In the following, whenever we write $F\rightarrow G$ for two filters $F$ and $G$, it will be understood that we are referring to this operation. Notice that for a frame $L$ and for $a,b\in L$ we have
${\uparrow}a\rightarrow{\uparrow}b=\{x\in L:b\leq x\vee a\}.$
This gives a useful characterization of exact filters.
Lemma 2.7 ([jakl24], Proposition 5.5).
A filter is exact if and only if it is the intersection of filters of the form ${\uparrow}a\rightarrow{\uparrow}b$ for some $a,b\in L$. In particular, if $F$ is an exact filter,
$F=\bigcap\{{\uparrow}a\rightarrow{\uparrow}b:b\leq a\vee f\mbox{ for all }f\in F\}.$
In particular, note that this means that for any $a\in L$ we have $\neg{\uparrow}a={\uparrow}a\to\{1\}=\{x\in L:x\vee a=1\}$. We say that a filter is regular if it is a regular element in the frame of filters (that is, if it is of the form $\neg F$ for some filter $F$). We call $\mathsf{Filt}_{\mathcal{R}}(L)$ the ordered collection of regular filters. Note that $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathsf{Filt}(L)$ is the Booleanization of the frame of $\mathsf{Filt}(L)$. Regular filters, too, have a useful concrete characterization.
Proposition 2.8 ([jakl24], Lemma 5.6).
The regular filters coincide with the intersections of filters of the form $\{x\in L:x\vee a=1\}$ for some $a\in L$. In particular, if $F$ is a regular filter,
$F=\{x\in L:x\vee f=1\mbox{ for all }f\in F\}.$
In the following, $\mathsf{Filt}_{\mathcal{CP}}(L)$ is the collection of completely prime filters and $\mathsf{Filt}_{\mathcal{SO}}(L)$ that of Scott-open filters, and $\mathcal{I}(-)$ denotes closure under set-theoretical intersections. Note that this includes the empty intersection, namely the whole frame $L$.
Theorem 2.9 ([jakl24], Corollary 5.11).
For any frame $L$, we have the following poset of sublocale inclusions:
[…]
Finally, we have the following characterizations of frame properties in terms of collections of filters.
Proposition 2.10 ([jakl24], Proposition 5.12).
For a frame $L$ we have:
  • $L$ is pre-spatial if and only if $\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$ contains all principal filters;
  • $L$ is spatial if and only if $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$ contains all principal filters;
  • $L$ is subfit if and only if $\mathsf{Filt}_{\mathcal{R}}(L)$ contains all principal filters.
2.3 Canonical extensions and the Prime Ideal Theorem
Canonical extensions
Canonical extensions for Boolean algebras were introduced by Jónsson and Tarski (see [jonsson51i] and [jonsson51ii]) in dealing with Boolean algebras with operators. The canonical extension may be seen as a pointfree version of the lattice of compact opens into the lattice of saturated sets of a Stone space. In fact, the Prime Ideal Theorem implies that the canonical extension of a Boolean algebra $B$ is the powerset of its dual space. Without assuming the Prime Ideal Theorem, the canonical extension of a Boolean algebra may be seen as the Booleanization of $\mathcal{U}(\mathsf{Filt}(B){\setminus}\{B\})$ – or, equivalently, the Booleanization of $\mathcal{U}(\mathsf{Idl}(B))$ – as shown in [bezhanishvili22]. Canonical extensions have also been introduced for distributive lattices. On this topic, we refer the reader to [gehrke94], [gehrke01], and [gehrke08]. If we assume the Prime Ideal Theorem, the canonical extension of any distributive lattice is isomorphic to the lattice of saturated sets of the corresponding coherent space. With no choice principles this is not necessarily the case.
In [jakl20], the question of what is the canonical extension of a frame is tackled for locally compact frames; there, the canonical extension of a general frame $L$ is defined as a monotone map $f^{\delta}:L\rightarrow L^{\delta}$ to a complete lattice $L^{\delta}$ such that the following two properties hold:
  1. 1.
    Density: every element of $L^{\delta}$ is a join of elements in $\{\bigwedge f[F]:F\in\mathsf{Filt}_{\mathcal{SO}}(L)\}$;
  2. 2.
    Compactness: for every Scott-open filter $F$ we have $\bigwedge f[F]\leq f(a)$ implies $a\in F$, for each $a\in L$.
Theorem 2.11.
([jakl20], Theorem 4.2) For a frame $L$, its canonical extension is unique, up to isomorphism. This is the map
$L\rightarrow\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op},$
$a\mapsto\bigcap\{F\in\mathsf{Filt}_{\mathcal{SO}}(L):a\in F\}.$
A frame $L$ is pre-spatial if whenever $a\nleq b$ there is a Scott-open filter containing $a$ and omitting $b$, for all $a,b\in L$.
Proposition 2.12.
([jakl20], Proposition 5.1) The map $f^{\delta}:L\rightarrow L^{\delta}$ is an injection if and only if $L$ is pre-spatial.
The Prime Ideal Theorem, pre-spatiality, and strict sobriety
In [erne07] the Prime Ideal Theorem – PIT hereon – is shown to be equivalent to the statement that every pre-spatial frame is also spatial. The so-called Strong Prime Element Theorem – which we will abbreviate as SPET – states that for every complete distributive lattice $D$, and any Scott-open filter $F\subseteq D$, for every element $a\in D$ not in $F$ there is a prime element $p\in D$ above $a$ with $p\notin F$. In [Banaschewski93] (Proposition 1) it is shown that PIT implies SPET. It is also known that SPET implies PIT. In [erne07] the notion of strict sobriety is introduced, and it is shown that sobriety implying strict sobriety is equivalent to the Ultrafilter Principle and several others choice principles. The concept was later developed in [erne18]. A space $X$ is strictly sober if it is $T_{0}$ and every Scott-open filter of its frame of opens is $\{U\in\Omega(X):F\subseteq U\}$ for some saturated set $F$, which is then necessarily compact. Strict sobriety is stronger than sobriety, even without assuming any choice principles.
3 Raney extensions
For a complete lattice $C$, we say that $L\subseteq C$ is a subframe of $C$ if $L$ equipped with the inherited order is a frame, and if the embedding $L\subseteq C$ preserves all joins and finite meets. A Raney extension is a pair $(L,C)$ such that $C$ is a coframe and $L$ is a subframe of $C$ such that:
  • The frame $L$ meet-generates $C$;
  • The embedding $L\subseteq C$ preserves strongly exact meets.
. We will sometimes use the expression Raney extension to refer to the coframe component of the pair, and for a pair $(L,C)$ we will say that this is a Raney extension of $L$, or that it is a Raney extension over $L$. A morphism of Raney extensions $f:(L,C)\rightarrow(M,D)$ is a coframe map $f:C\rightarrow D$ such that whenever $a\in L$ we have $f(a)\in M$ and such that the restriction $f|_{L}:L\rightarrow M$ is a frame map. We call $\mathbf{Raney}$ the category of Raney extensions with Raney maps.
Example 3.1.
For a topological space $X$, the pair $(\Omega(X),\mathcal{U}(X))$ is a Raney extension. That strongly exact meets are preserved by the embedding $\Omega(X)\subseteq\mathcal{U}(X)$ is the content of Proposition 5.3 of [ball14].
We shall see more concrete examples of Raney extensions in Subsection 3.1. Observe that for any Raney extension $(L,C)$, by the universal property of the ideal completion of a distributive lattice, there is a coframe surjection $\mathsf{Filt}(L)^{op}\to C$. The fact that it extends the map $L\subseteq C$, and that it preserves all meets, means that this map is defined as $F\mapsto\bigwedge F$ for each filter $F\in\mathsf{Filt}(L)$. As it is a coframe map, it has a left adjoint. Now, notice that for each filter $F\in\mathsf{Filt}(L)$ and each $c\in C$ we have:
$c\leq\bigwedge F\mbox{ if and only if }c\leq f\mbox{ for all $f\in F$, if and only if $F\subseteq{\uparrow}c\cap L$}.$
The equivalences above mean that the left adjoint we are looking for acts as $c\mapsto{\uparrow}c\cap L$, and from now on we will denote this map as ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$. The fixpoints on the coframe of filters are then exactly those of the form ${\uparrow}^{L}c$ for some $c\in C$. As the starting map $\bigwedge$ is a coframe map, these form a subcolocale, which we will hereon call $C^{*}\subseteq\mathsf{Filt}(L)^{op}$. On the other hand, all elements of $C$ are fixpoints, as $\bigwedge:\mathsf{Filt}(L)^{op}\to C$ is a surjection. Let us state these observations as a theorem.
Theorem 3.1.
For a Raney extension $(L,C)$, we have an adjunction
$\bigwedge:\mathsf{Filt}(L)^{op}\leftrightarrows C:{\uparrow}^{L},$
which maximally restricts to a pair of mutually inverse isomorphisms
$\bigwedge:C^{*}\leftrightarrows C:{\uparrow}^{L}.$
These are also isomorphisms of Raney extensions $\bigwedge:(L,C^{*})\leftrightarrows(L,C):{\uparrow}^{L}$.
Proof.
It only remains to show that the isomorphisms ${\uparrow}^{L}:C\leftrightarrows C^{*}:\bigwedge$ restrict correctly to the frame components. We notice that for $a\in L$, indeed, the filter ${\uparrow}^{L}a$ is the principal filter ${\uparrow}a\subseteq L$, an element of the generating frame of $C^{*}$. Conversely, any principal filter of $L$ is of this form. ∎
Let us now tie the notion of Raney extension with that of canonical extension. For a monotone map $f:L\to C$ of a lattice $L$ into a complete lattice $C$, we introduce the following two properties:
  1. 1.
    $\mathcal{F}$-density: the collection $\{\bigwedge f[F]:F\in\mathcal{F}\}$ join-generates $C$;
  2. 2.
    $\mathcal{F}$-compactness: whenever $\bigwedge f[F]\leq f(a)$ we also have $a\in F$ for every $F\in\mathcal{F}$ and every $a\in L$.
We say that the map is $\mathcal{F}$-canonical if and only if it is both $\mathcal{F}$-dense and $\mathcal{F}$-compact. For brevity, in the following we will refer to $\mathsf{Filt}_{\mathcal{SO}}(L)$-canonicity simply as $\mathcal{SO}$-canonicity, and analogously for all other similarly denoted collections of filters, and for density and compactness.
Example 3.2.
For a sober space $X$, the pair $(\Omega(X),\mathcal{U}(X))$ is a $\mathcal{SO}$-canonical Raney extension, provided that the Prime Ideal Theorem holds. This is observed in Example 3.5 of [jakl20]. On the one hand, the coframe $\mathcal{U}(X)$ is join-generated by elements of the form ${\uparrow}x$ for $x\in X$, and these are intersections of neighborhood filters, which are completely prime, hence Scott-open. Then, the extension is $\mathcal{SO}$-dense. For $\mathcal{SO}$-compactness, we rely on the Hofmann-Mislove Theorem. If $F\subseteq\Omega(X)$ is a Scott-open filter, then by the Theorem it must be $\{U\in\Omega(X):\bigcap F\subseteq U\}$, and so, indeed, for every open $U$ we have $\bigcap F\subseteq U$ implies $U\in F$. Recall that the Hofmann-Mislove Theorem is dependent on the Prime Ideal Theorem, see for instance [erne18], Theorem 3. With Proposition 3.11 we will see that if we replace Scott-open by completely prime, we have an analogous result which does not rely on the Prime Ideal Theorem. We will explore the relation between the Prime Ideal Theorem and $\mathcal{SO}$-canonicity in Section 5.
Theorem 3.1 tells us that for a Raney extension $(L,C)$ we may identify elements of $C$ with filters of $L$. This proves to be useful when dealing with density and compactness. In the following, for a collection of filters $\mathcal{F}$, we denote as $\mathcal{F}^{*}$ the collection $\{{\uparrow}^{L}\bigwedge F:F\in\mathcal{F}\}$.
Proposition 3.2.
For any Raney extension $(L,C)$ and any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$,
  1. 1.
    $(L,C)$ is $\mathcal{F}$-dense if and only if $C^{*}\subseteq\mathcal{I}(\mathcal{F}^{*})$;
  2. 2.
    $(L,C)$ is $\mathcal{F}$-compact if and only if $\mathcal{F}\subseteq C^{*}$.
In particular, $(L,C)$ is $\mathcal{F}$-canonical if and only if $\mathcal{I}(\mathcal{F})^{op}=C^{*}$.
Proof.
Let us prove the first claim. If $(L,C)$ is $\mathcal{F}$-dense, then for all $c\in C$ we have that $c=\bigvee_{i}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$, that is, ${\uparrow}^{L}c={\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}$. As ${\uparrow}^{L}:C\to\mathsf{Filt}(L)^{op}$ is a left adjoint, it preserves all joins, and so ${\uparrow}^{L}\bigvee_{i}\bigwedge F_{i}=\bigcap_{i}{\uparrow}^{L}\bigwedge F$. For the converse, suppose that $C^{*}\subseteq\mathcal{I}(\mathcal{F}^{*})$. We have that, for $c\in C$, ${\uparrow}^{L}c=\bigcap_{i}{\uparrow}^{L}\bigwedge F_{i}$ for some collection $F_{i}\in\mathcal{F}$. Again, by preservation of joins of ${\uparrow}^{L}$, we obtain $c=\bigvee_{i}\bigwedge F_{i}$. To see the equivalence stated in the second claim, we observe that for any filter $F\subseteq L$ we always have $F\subseteq{\uparrow}^{L}\bigwedge F$. For any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ it is the case that for all $F\in\mathcal{F}$ we have the reverse set inclusion if and only if the Raney extension is $\mathcal{F}$-compact. But this is also equivalent to having that all filters in $\mathcal{F}$ are fixpoints of ${\uparrow}^{L}\dashv\bigwedge$, i.e. them being elements of $C^{*}$. ∎
Corollary 3.3.
For any collection of filters $\mathcal{F}\subseteq\mathsf{Filt}(L)$, a Raney extension $(L,C)$ such that $C^{*}\subseteq\mathcal{I}(\mathcal{F}\cap C^{*})$ is $\mathcal{F}$-dense.
Existence and uniqueness of $\mathcal{F}$-canonical extensions of lattices to complete lattices are well-known, and these results stem from the theory of polarities by Birkhoff (see [birkhoff40]). For a general version of the existence and uniqueness results, see for instance Section 2 of [gehrke06], see [gehrke01] for its application to distributive lattices. From particularizing the analysis of [gehrke01] to the case where we start from a frame, we directly obtain the following.
Theorem 3.4.
(see for example [gehrke01], in particular Remark 2.8) For a frame $L$ and a collection $\mathcal{F}\subseteq L$ of its filters such that $\mathcal{I}(\mathcal{F})$ contains the principal ones, there is a unique injective monotone map $f^{\mathcal{F}}:L\rightarrow L^{\mathcal{F}}$ to a complete lattice $L^{\mathcal{F}}$ which is $\mathcal{F}$-canonical. Concretely, this is the embedding $L\subseteq\mathcal{I}(\mathcal{F})^{op}$ mapping each element to its principal filter. This embedding also preserves the frame operations, and $L$ meet-generates $\mathcal{I}(\mathcal{F})^{op}$.
We now wish to adapt the theorem above to prove existence and uniqueness of $\mathcal{F}$-canonical Raney extensions on a frame $L$. An extension as required will not exist for all choices of $\mathcal{F}$, because for some choices of $\mathcal{F}$ the structure $(L,\mathcal{I}(\mathcal{F})^{op})$ will not be a Raney extension.
Lemma 3.5.
Suppose that $(L,C)$ is a Raney extension. All the following hold.
  • $C^{*}$ contains all principal filters;
  • $C^{*}\subseteq\mathsf{Filt}(L)^{op}$ is a subcolocale inclusion;
  • All filters in $C^{*}$ are strongly exact.
Proof.
For the first item, we only notice that it is clear that $a=\bigwedge{\uparrow}^{L}a$ for all $a\in L$. The second item is one of the observations before Theorem 3.1. Let us show the third item. Suppose that $F\in C^{*}$, and that $x_{i}\in F$ is a family such that the meet $\bigwedge^{L}_{i}x_{i}$, as calculated in $L$, is strongly exact. By definition of Raney extension, this meet is preserved by the embedding $e:L\subseteq C$. This means that $\bigwedge^{L}_{i}x_{i}=\bigwedge_{i}x_{i}$, where the second meet is computed in $C$. Therefore, since $\bigwedge F\leq x_{i}$ for all $i\in I$, we also have $\bigwedge F\leq\bigwedge_{i}x_{i}$. Since $F\in C^{*}$, we have $F={\uparrow}^{L}\bigwedge F$, and so $\bigwedge^{L}_{i}x_{i}\in F$. ∎
Theorem 3.6.
For a frame $L$ and any collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ of filters, the $\mathcal{F}$-canonical Raney extension exists if and only if:
  1. 1.
    $\mathcal{I}(\mathcal{F})$ contains all principal filters;
  2. 2.
    $\mathcal{I}(\mathcal{F})^{op}\subseteq\mathsf{Filt}(L)^{op}$ is a subcolocale inclusion;
  3. 3.
    All filters in $\mathcal{F}$ are strongly exact.
In case it exists, it is unique, up to isomorphism. Concretely, it is described as the structure coming from the theory of polarities of Birkhoff, that is, the pair $(L,\mathcal{I}(\mathcal{F})^{op})$.
Proof.
Let us show that the three conditions are necessary. By Proposition 3.2, if an $\mathcal{F}$- canonical extension $(L,C)$ exists then $C^{*}=\mathcal{I}(\mathcal{F})^{op}$. Necessity then follows by Lemma 3.5. Let us now show that for a collection $\mathcal{F}\subseteq\mathsf{Filt}(L)$ satisfying the three properties above, the pair $(L,\mathcal{I}(\mathcal{F})^{op})$ is a Raney extension. We know from Theorem 3.4 that $L\subseteq\mathcal{I}(\mathcal{F})^{op}$ preserves the frame operations (and this is also easy to check), and that $L$ meet-generates the coframe component. We show that the embedding $L\subseteq\mathcal{I}(\mathcal{F})^{op}$ preserves strongly exact meets. Suppose that $x_{i}\in L$ is a family such that their meet $\bigwedge^{L}_{i}x_{i}$ is strongly exact. As all filters in $\mathcal{I}(\mathcal{F})$ are strongly exact, any such filter which contains ${\uparrow}x_{i}$ for all $i\in I$ must also contain $\bigwedge^{L}_{i}x_{i}$. This means that in the coframe $\mathcal{I}(\mathcal{F})^{op}$ the greatest lower bound of the family $\{{\uparrow}x_{i}:i\in I\}$ is the principal filter ${\uparrow}\bigwedge^{L}_{i}x_{i}$. This means that the meet $\bigwedge^{L}_{i}x_{i}$ is preserved. The fact that it satisfies the required universal property follows from the characterization in Proposition 3.2, and uniqueness follows from Theorem 3.4. ∎
3.1 Notable examples of Raney extensions
In this subsection, we look at some concrete examples of Raney extensions.
Proposition 3.7.
The following are all Raney extensions.
  • The pair $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ for any frame $L$;
  • The pair $(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ for any frame $L$;
  • The pair $(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op})$ for subfit $L$;
  • The pair $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$ for pre-spatial $L$;
  • The pair $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ for spatial $L$.
Proof.
We use the characterization in Theorem 3.6. That the collections of filters below are subcolocales of $\mathsf{Filt}_{\mathcal{SE}}(L)^{op}$ follows from Theorem 2.9. The collections of exact and of strongly exact filters contain the principal filters for every frame, whereas for the last three items we refer to Proposition 2.10. ∎
Because of the isomorphisms in Theorem 2.6, for any frame $L$ the following embeddings into coframes are Raney extensions, up to isomorphism.
  • $\mathfrak{o}:L\to\mathsf{S}_{\mathfrak{o}}(L)$,
  • $\mathfrak{c}:L\to\mathsf{S}_{\mathfrak{c}}(L)^{op}$.
We look at Raney extensions in relation to the concrete lattice of saturated sets of a space. We now expand a result in [jakl24] (Lemma 5.4) which states an isomorphism $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\cong\mathcal{U}(\mathsf{pt}(L))$, and show that it is the restriction of an adjunction which may be seen as the point-set version of that of Theorem 2.6. In the following $\varphi$ is the spatialization map of the frame $L$.
Theorem 3.8.
For a frame $L$, there is an adjunction
[…]
$F\mapsto\bigcap\varphi[F],$
[…]
Proof.
[…]
Corollary 3.9.
For any spatial frame $L$, the pair $(L,\mathcal{U}(\mathsf{pt}(L)))$ is a Raney extension, where we have identified each element $a\in L$ with the open $\varphi(a)$.
Proof.
For a spatial frame $L$, the pair $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op})$ is a Raney extension. By Theorem 3.8, there is an isomorphism $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\cong\mathcal{U}(\mathsf{pt}(L))$, and by definition of the isomorphism this maps principal filters to opens of $\mathsf{pt}(L)$. ∎
Let us now consider the motivating example behind the definition of Raney extension.
Lemma 3.10.
For every topological space $X$ the Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-dense.
Proof.
Let $X$ be a topological space. Consider some family $U_{i}\subseteq X$ of open sets. To show the claim, we use Corollary 3.3, and show that ${\uparrow}^{\Omega(X)}\bigcap_{i}U_{i}$ is an intersection of completely prime filters in $\mathcal{U}(X)^{*}$. The filter may be written as
$\bigcap\{N(x):x\in\bigcap_{i}U_{i}\},$
where $N(x)$ denotes the neighborhood filter of a point $x\in X$. Indeed, a neighborhood filter $N(x)$ is $\{U\in\Omega(X):{\uparrow}x\subseteq U\}$, and so it is in $\mathcal{U}(X)^{*}$. ∎
Proposition 3.11.
A $T_{0}$ topological space $X$ is sober if and only if $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact. In particular, a $T_{0}$ space $X$ is sober if and only if the pair $(\Omega(X),\mathcal{U}(X))$ is a realization of the $\mathcal{CP}$-canonical Raney extension.
Proof.
Suppose that $X$ is a sober space, and let $P$ be a completely prime filter. By definition of $\mathcal{CP}$-compactness, we have to show that $\bigcap P\subseteq U$ implies that $U\in P$ for every open $U\subseteq X$. By sobriety, each completely prime filter is the neighborhood filter of some point, and so we may assume $P=N(x)$ for some $x\in X$. We have $\bigcap N(x)={\uparrow}x$, and so, indeed, $\bigcap N(x)\subseteq U$ implies that $x\in U$, and so $U\in N(x)$. For the second part of the claim, assume that for some $T_{0}$ space $X$ the pair $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact. Let us show that in $X$ all completely prime filters are neighborhood filters. Let $P\subseteq\Omega(X)$ be a completely prime filter. By the characterization in Proposition 3.2, this is of the form $\{U\in\Omega(X):Y\subseteq U\}$, for some saturated set $Y$. By complete primality, the set $Y$ is then completely join-prime in the lattice $Y$, and so it is a principal upset ${\uparrow}x$ for some $x\in X$. Hence $P=N(x)$. This means that $X$ is sober. The last part of the claim follows by combining this characterization of sobriety with Lemma 3.10. ∎
4 Extensions of frame maps
We ask when a map $f:L\to M$ of frames can be extended to Raney extensions of these frames. For a collection of filters $\mathcal{F}$ and an arbitrary filter $G$ of a frame $L$, we introduce the following closure operator in $\mathsf{Filt}(L)$:
$\mbox{cl}_{\mathcal{F}}(G)=\bigcap\{F\in\mathcal{F}:G\subseteq F\}.$
In case $\mathcal{F}=\mathsf{Filt}_{\mathcal{SO}}(L)$, as an abbreviation we will denote this closure operator as $\mbox{cl}_{\mathcal{SO}}$, and similarly for similarly named collections of filters. We note that for a subcolocale $\mathcal{F}\subseteq\mathsf{Filt}(L)^{op}$, the meet of a family $F_{i}$ is computed as $\mbox{cl}_{\mathcal{F}}(\bigcup_{i}F_{i})$. Recall that, by Theorem 3.1, we may identify elements of the coframe components of Raney extensions with filters of the form ${\uparrow}^{L}a$. Then, if we are seeking to extend a frame map $f:L\to M$ to Raney extensions $(L,C)$ and $(M,D)$, we are looking for a coframe morphism making the following commute:
[…]
By commutativity of the diagram, if this map exists, it has to act as ${\uparrow}a\mapsto{\uparrow}f(a)$ for each $a\in L$. Because it also has to be meet-preserving, and because the order on $C^{*}$ and $D^{*}$ is reverse set inclusion, for a filter $F\in C^{*}$ it has to act as $\bigcup\{{\uparrow}f:f\in F\}\mapsto\mbox{cl}_{D^{*}}(\bigcup\{{\uparrow}f(x):x\in F\})$. This means that if such lifting exists, then it has to be defined as $F\mapsto\mbox{cl}_{D^{*}}(f[F])$.
Lemma 4.1.
Suppose that we have coframes $C$ and $D$, such that $L\subseteq C$ is a subframe that meet-generates $C$. Suppose that a map $f:C\rightarrow D$ preserves all meets, as well as all the joins of $L$. Then it is a coframe map.
Proof.
To show that it is a coframe map, it suffices to show that it preserves finite joins. Consider $x,y\in C$. We have $x=\bigwedge_{i}x_{i}$ and $y=\bigwedge_{j}y_{j}$ for families $x_{i},y_{j}\in L$. Then
$f(x\vee y)=f(\bigwedge_{i}x_{i}\vee\bigwedge_{j}y_{j})=\bigwedge_{i,j}f(x_{i})\vee f(y_{j})=f(x)\vee f(y).\qed$
Lemma 4.2.
Suppose that we have a frame map $f:L\rightarrow M$ and Raney extensions $(L,C)$ and $(M,D)$. If the map $\mathsf{cl}_{D^{*}}(f[-]):C^{*}\to D^{*}$ preserves all meets, then it is a coframe map.
Proof.
It suffices to observe that we are in the situation described by Lemma 4.1. This is because the map $\mbox{cl}_{D^{*}}(f[-])$ lifts a frame map $f:L\rightarrow M$, and so its restriction to $L$ is the composition of the join-preserving map $f$ with the join-preserving inclusion $M\subseteq D^{*}$. ∎
For a poset map to preserve all meets, it suffices for it to have a left adjoint. In general, for all $F\in C^{*}$ and $G\in D^{*}$, we have $\mbox{cl}_{D^{*}}(f[F])\subseteq G$ if and only if $f[F]\subseteq G$, and this holds if and only if $F\subseteq f^{-1}(G)$. As the order on $C^{*}$ and $D^{*}$ is reverse set inclusion, if a well-defined preimage map $f^{-1}:D^{*}\to C^{*}$ exists, then it is the left adjoint we are looking for. A frame map $f:L\to M$ always determines a preimage (frame) map $f^{-1}:\mathsf{Filt}(M)\to\mathsf{Filt}(L)$. The issue with the preimage is that it is not guaranteed to map filters of $D^{*}$ to filters of $C^{*}$.
Theorem 4.3.
Suppose that $f:L\rightarrow M$ is a frame map and that we have Raney extensions $(L,C)$ and $(M,D)$. The map $f$ can be extended to a map of these Raney extensions if and only if $f^{-1}(G)\in C^{*}$ for every $G\in D^{*}$.
Proof.
It remains to show that, if a left adjoint of $\mbox{cl}_{D^{*}}(f[-])$ exists, then it has to be the preimage map. For a left adjoint $\mathcal{L}$ and for each $F\in C^{*}$ and $G\in D^{*}$, by definition of the closure $\mbox{cl}_{D^{*}}$, we would have $F\subseteq\mathcal{L}(G)$ if and only if $f[F]\subseteq G$. For each $x\in L$, we have $x\in f^{-1}(G)$ if and only if $f(x)\in G$, and this is in turn equivalent to $f[{\uparrow}x]\subseteq G$, as noticed above this holds if and only if ${\uparrow}x\subseteq\mathcal{L}(G)$, that is $x\in\mathcal{L}(G)$. For the last step, we have used the fact that ${\uparrow}x\in C^{*}$, as this collection is assumed to contain all principal filters of $L$. We have then shown that, for each $x\in L$, and each $G\in D^{*}$, $x\in f^{-1}(G)$ if and only if $x\in\mathcal{L}(G)$. ∎
We also notice that for a frame map $f:L\to M$ which can be extended to a Raney map $(L,C)\to(M,D)$, the definition of the extension as a coframe map $C\to D$ has to be $c\mapsto\bigwedge\{f(a):a\in{\uparrow}^{L}c\}$, from the fact that it has to extend $f$, together with preservation of all meets. We apply Theorem 4.3 above to some of the Raney extensions that we have seen.
Lemma 4.4.
Any frame morphism $f:L\rightarrow M$ lifts to a coframe morphism
$f_{\mathcal{SE}}:\mathsf{Filt}_{\mathcal{SE}}(L)^{op}\rightarrow\mathsf{Filt}_{\mathcal{SE}}(M)^{op}.$
Proof.
By Theorem 4.3, a frame morphism $f:L\rightarrow M$ lifts as required if preimages of strongly exact filters are strongly exact. Suppose, then, that $F\subseteq M$ is strongly exact. Suppose that the meet $\bigwedge_{i}x_{i}$ is strongly exact, and that $f(x_{i})\in F$. Because all frame morphisms preserve strongly exact meets, as well as strong exactness of meets, we have $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})\in F$, as desired. ∎
Corollary 4.5.
Every map of frames $L\to M$ lifts to a coframe map $\mathsf{S}_{\mathfrak{o}}(L)\to\mathsf{S}_{\mathfrak{o}}(M)$.
It is known that, for all frame maps, preimages of completely prime filters are completely prime. Recall that the preimage map also preserves arbitrary intersections. This, together with Theorem 4.3, gives us the following.
Lemma 4.6.
Any frame morphism $f:L\rightarrow M$ between spatial frames lifts to a coframe morphism
$f_{\mathcal{CP}}:\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))^{op}\rightarrow\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(M))^{op}.$
Lemma 4.7.
Any frame morphism $f:L\rightarrow M$ between pre-spatial frames lifts to a coframe morphism
$f_{\mathcal{SO}}:\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op}\rightarrow\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}.$
Proof.
By Theorem 4.3, it suffices to show that for a frame morphism $f:L\rightarrow M$ between pre-spatial frames preimages of Scott-open filters are Scott-open. Suppose that $F\subseteq L$ is a Scott-open filter, and that $\{x_{i}:i\in I\}\subseteq L$ is a directed family such that $f(\bigvee_{i}x_{i})=\bigvee f(x_{i})\in F$. Observe that the family $\{f(x_{i}):i\in I\}$ is directed, and so by Scott-openness of $F$ we must have $f(x_{i})\in F$ for some $i\in I$, as desired. ∎
There is a forgetful functor $\pi_{1}:\mathbf{Raney}\rightarrow\mathbf{Frm}$ which forgets about the second component of the extension. We observe that the assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ on objects can be extended to a functor $\mathsf{Filt}_{\mathcal{SE}}$ by mapping each frame morphism to the morphism $f_{\mathcal{SE}}$ whose existence is established by Lemma 4.4.
Theorem 4.8.
For a frame $L$, the pair $(L,\mathsf{Filt}_{\mathcal{SE}}(L)^{op})$ is the free Raney extension over it, that is, we have $\mathsf{Filt}_{\mathcal{SE}}\dashv\pi_{1}$. In particular, the category of frames is a full coreflective subcategory of $\mathbf{Raney}$.
Proof.
Suppose that we have a frame map $f:L\rightarrow M$. Let $(M,D)$ be a Raney extension. By Lemma 3.5, we have $D^{*}\subseteq\mathsf{Filt}_{\mathcal{SE}}(M)^{op}$, and by Lemma 4.4, preimages of strongly exact filters are strongly exact. Therefore, preimages of elements in $D^{*}$ are in $\mathsf{Filt}_{\mathcal{SE}}(L)$. By Theorem 4.3, the frame map lifts as required, and so we have $\mathsf{Filt}_{\mathcal{SE}}\dashv\pi_{1}$. ∎
We use the isomorphism in Theorem 2.6 to prove a universal property of the collection $\mathsf{S}_{\mathfrak{o}}(L)$ of fitted sublocales of a frame.
Corollary 4.9.
For a frame $L$, the pair $(L,\mathsf{S}_{\mathfrak{o}}(L))$ is the free Raney extension on it.
We already have seen that each frame has the largest Raney extension. Every frame also has the smallest Raney extension.
Lemma 4.10.
For a frame $L$ the collection $\mathsf{Filt}_{\mathcal{E}}(L)$ is the smallest sublocale of $\mathsf{Filt}(L)$ containing all the principal filters.
Proof.
Let $\mathcal{S}\subseteq\mathsf{Filt}(L)$ be a sublocale containing all the principal filters. For any $x,y\in L$, we must have ${\uparrow}x\rightarrow{\uparrow}y\in\mathcal{S}$. As sublocales are closed under all meets, all intersections of filters of the form ${\uparrow}x\rightarrow{\uparrow}y$ must be in $\mathcal{S}$. Therefore, by the characterization in Lemma 2.7, $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathcal{S}$. ∎
We provide the frame version of a result in [ball14b]: in Theorem 3.7, it is shown that for a meet-semilattice $S$ the smallest frame generated by it is $\mathcal{J}^{e}(S)$, the collection of all downsets which are closed under those joins of $S$ that distribute over all finite meets. Recently, the same result has been re-proven for frames with bases of meet-semilattices in [bezhanishvili24].
Proposition 4.11.
For all Raney extensions $(L,C)$ we have a surjection
$(L,C)\rightarrow(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op}),$
$c\mapsto\mbox{cl}_{\mathcal{E}}({\uparrow}^{L}c).$
Proof.
The collection $C^{*}\subseteq\mathsf{Filt}(L)$ is a sublocale which contains all principal filters, and so $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$ by Lemma 4.10. This can be rephrased as saying that the preimage map relative to the identity $1_{L}:L\to L$, which is itself the identity on $\mathsf{Filt}(L)$, maps exact filters to filters of $C^{*}$. Then by Theorem 4.3 we have a surjection $c\mapsto\mbox{cl}_{\mathcal{E}}({\uparrow}^{L}c)$. ∎
Corollary 4.12.
For a frame $L$, the smallest Raney extension on it is $(L,\mathsf{S}_{\mathfrak{c}}(L)^{op})$.
We may order Raney extensions over some frame $L$ by subcolocale inclusion of the coframe components. We obtain a result which may be seen as a version for Raney extensions of Theorem 3.7 in [ball14ext], where the authors consider the ordered collections of all frames join-generated by a distributive lattice. In the recent work [bezhanishvili24] the result is given a new proof.
Theorem 4.13.
For a frame $L$, the ordered collection of Raney extensions over $L$ is the interval $[\mathsf{Filt}_{\mathcal{E}}(L),\mathsf{Filt}_{\mathcal{SE}}(L)]$ of the coframe of sublocales of $\mathsf{Filt}_{\mathcal{SE}}(L)$.
Proof.
That every Raney extension belongs to the section $[\mathsf{Filt}_{\mathcal{E}}(L),\mathsf{Filt}_{\mathcal{SE}}(L)]$ follows from Theorem 4.8 and Proposition 4.11. Suppose that we have a sublocale $\mathcal{F}\subseteq\mathsf{Filt}_{\mathcal{SE}}(L)$ such that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq\mathcal{F}$. By Lemma 4.10, $\mathcal{F}$ contains all principal filters, and so, by Theorem 3.6, the pair $(L,\mathcal{F}^{op})$ is a Raney extension. ∎
4.1 Functoriality of the exact filters extension
The assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$ is not always functorial, as shown in [ball19], where the question is explored for structure $\mathsf{S}_{\mathfrak{c}}(L)$, isomorphic to $\mathsf{Filt}_{\mathcal{E}}(L)$. For a frame morphism $f:L\rightarrow M$, we will say that it is exact if whenever the meet of a family $\{x_{i}:i\in I\}\subseteq L$ is exact, so is the meet of $\{f(x_{i}):i\in I\}$, and furthermore $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$.
Proposition 4.14.
A morphism $f:L\rightarrow M$ is exact if and only if preimages of exact filters are exact. This holds if and only if the morphism can be extended to a morphism
$f_{\mathcal{E}}:(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})\rightarrow(M,\mathsf{Filt}_{\mathcal{E}}(M)^{op}).$
Proof.
Suppose that $f:L\rightarrow M$ is an exact frame map, and that $G\subseteq M$ is an exact filter. Suppose that $\bigwedge_{i}x_{i}\in L$ is an exact meet such that $f(x_{i})\in G$. By exactness of this map, the meet $\bigwedge_{i}f(x_{i})$ is exact and so $\bigwedge_{i}f(x_{i})\in G$. Again, by exactness of $f$, we have $\bigwedge_{i}f(x_{i})=f(\bigwedge_{i}x_{i})$. Indeed, then, $\bigwedge_{i}x_{i}\in f^{-1}(G)$. Conversely, suppose that we have a frame map $f:L\rightarrow M$ such that it is not exact. This means that either we have an exact meet $\bigwedge_{i}x_{i}\in L$ such that it is not preserved by $f$, or we have an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. We consider these two cases in turn. In the first case, we consider the principal filter ${\uparrow}\bigwedge_{i}f(x_{i})$. This is exact, as it is closed under all meets. We notice that by our hypothesis $f(\bigwedge_{i}x_{i})$ is not an element of this filter. Let us call $F$ the preimage of this filter. We have that $x_{i}\in F$ but $\bigwedge_{i}x_{i}\notin F$, and so $F$ is not exact. In the second case, consider an exact meet $\bigwedge_{i}x_{i}\in L$ such that $\bigwedge_{i}f(x_{i})$ is not exact. In particular, let $y\in M$ be such that $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$. We now consider the exact filter
${\uparrow}y\rightarrow{\uparrow}\bigwedge_{i}(f(x_{i})\vee y)=\{m\in M:\bigwedge_{i}(f(x_{i})\vee y)\leq y\vee m\}.$
That this is an exact filter follows from the characterization of Lemma 2.7. Let $F$ be the preimage of this filter. We have that $x_{i}\in F$ for each $i\in I$. We claim that $\bigwedge_{i}x_{i}\notin F$. This follows from the fact that by our hypothesis $\bigwedge_{i}(f(x_{i})\vee y)\nleq(\bigwedge_{i}f(x_{i}))\vee y$ and $f(\bigwedge_{i}x_{i})\leq\bigwedge_{i}f(x_{i})$. The rest of the claim follows by Theorem 4.3. ∎
Let us call $\mathbf{Frm}_{\mathcal{E}}$ the category of frames with exact maps. Let us also call $\mathbf{Raney}_{\mathcal{E}}$ the category of Raney extensions with morphisms that, restricted to the frame components, are exact. The assignment $L\mapsto(L,\mathsf{Filt}_{\mathcal{E}}(L)^{op})$, then, determines a functor $\mathsf{Filt}_{\mathcal{E}}:\mathbf{Frm}_{\mathcal{E}}\to\mathbf{Raney}_{\mathcal{E}}$.
Theorem 4.15.
We have an adjunction $\pi_{1}:\mathbf{Raney}_{\mathcal{E}}\leftrightarrows\mathbf{Frm}_{\mathcal{E}}:\mathsf{Filt}_{\mathcal{E}}$ with $\pi_{1}\dashv\mathsf{Filt}_{\mathcal{E}}$.
Proof.
Suppose that we have an exact frame map $f:L\rightarrow M$, and that $(L,C)$ is a Raney extension. By Proposition 4.14, as $f$ is exact, preimages of filters in $\mathsf{Filt}_{\mathcal{E}}(M)$ are in $\mathsf{Filt}_{\mathcal{E}}(L)$. Furthermore, $C^{*}$ must contain all principal filters, and so by Lemma 4.10 this implies that $\mathsf{Filt}_{\mathcal{E}}(L)\subseteq C^{*}$. Then, preimages of exact filters of $M$ are in $C^{*}$. By Theorem 4.3, then, we have a map of Raney extensions $(L,C)\rightarrow(M,\mathsf{Filt}_{\mathcal{E}}(M)^{op})$ extending $f$, as desired. ∎
5 Canonical extensions as Raney extensions
We now look at the notion of canonical extension of a frame from [jakl20], and see how it relates to Raney extensions. We will also look at how the Prime Ideal Theorem relates to some of our results. For a pre-spatial frame $L$, we will call its canonical extension the Raney extension
$(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op}).$
This is indeed a Raney extension, by Proposition 2.10. For a Raney extension $(L,C)$, we say that an element $c\in C$ is compact if, for every directed collection $D\subseteq L$, we have that $c\leq\bigvee D$ implies that $c\leq d$ for some $d\in D$. We say that a Raney extension $(L,C)$ is algebraic if every element of $c$ is the join of compact elements.
Lemma 5.1.
A Raney extension $(L,C)$ is algebraic if and only if $C^{*}\subseteq\mathcal{I}(C^{*}\cap\mathsf{Filt}_{\mathcal{SO}}(L))$.
Proof.
Notice that an element $x\in C$ is compact if and only if the filter ${\uparrow}^{L}x$ is Scott-open. Consider the isomorphism ${\uparrow}^{L}:C\cong C^{*}$. The Raney extension $(L,C)$ is algebraic if and only if in $C^{*}$ every element is a join of Scott-open filters of the form ${\uparrow}^{L}x$ for some $x\in L$. We have that $C^{*}\subseteq\mathsf{Filt}(L)^{op}$ is a subcolocale inclusion, and subcolocale inclusions preserves all joins, and joins in $\mathsf{Filt}(L)^{op}$ are intersections. Therefore, algebraicity of $(L,C)$ is equivalent to every filter in $C^{*}$ being an intersection of Scott-open filters in $C^{*}$. ∎
Lemma 5.2.
A frame admits an algebraic Raney extension if and only if it is pre-spatial.
Proof.
First, we observe that if a frame admits an algebraic Raney extension this means that principal filters must all be intersections of Scott-open filters, by Lemma 5.1. By Proposition 2.10, the frames with this property are exactly the pre-spatial ones. For a pre-spatial frame $L$, an algebraic Raney extension is $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$. ∎
We are now ready to characterize canonical extensions of frames as free algebraic Raney extensions.
Theorem 5.3.
For a pre-spatial frame $L$, its canonical extension is the free algebraic Raney extension over it. That is, whenever we have a frame map $f:L\rightarrow M$ and $(M,C)$ is an algebraic Raney extension, the map $f$ can be extended to a map of Raney extensions
$(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,C).$
Proof.
The assignment is functorial, by Lemma 4.7. Suppose that $L$ is a pre-spatial frame, and that $(M,C)$ is an algebraic Raney extension. Suppose that we have a frame map $f:L\rightarrow M$. Consider the canonical extension $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})$. We have that, as $(M,C)$ is algebraic, $C^{*}\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(M))^{op}$, by Lemma 5.1. By Lemma 4.7, preimages of Scott-open filters are Scott-open. Then, preimages of filters in $C^{*}$ are in $\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$. By Theorem 4.3 this means that there is a map of Raney extensions $(L,\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))^{op})\rightarrow(M,C)$ extending the frame map $f:L\rightarrow M$. ∎
We now look at the role of the Prime Ideal Theorem, and work towards characterizing sobriety and strict sobriety of spaces.
Lemma 5.4.
A $T_{0}$ space $X$ is strictly sober if and only if $(\Omega(X),\mathcal{U}(X))$ is a $\mathcal{SO}$-compact Raney extension.
Proof.
By Proposition 3.2, a Raney extension is $\mathcal{SO}$-compact if and only if $\mathsf{Filt}_{\mathcal{SO}}(L)\subseteq C^{*}$. The claim follows by definition of strict sobriety. ∎
Proposition 5.5.
A $T_{0}$ space is strictly sober if and only if $(\Omega(X),\mathcal{U}(X))$ is the canonical extension of $\Omega(X)$.
Proof.
If $X$ is strictly sober, $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{SO}$-compact, by Lemma 5.4. It is also $\mathcal{CP}$-dense, by Lemma 3.10, and as completely prime filters are Scott-open it is also $\mathcal{SO}$-dense. Conversely, if $X$ is a space such that $(\Omega(X),\mathcal{U}(X))$ is a canonical extension, in particular this Raney extension is $\mathcal{SO}$-compact, hence $X$ is strictly sober by Lemma 5.4. ∎
It is known that if we do not assume choice principles it is not the case that sobriety implies strict sobriety. Let us look at a counterexample for this.
Example 5.1.
We assume the negation of the Ultrafilter Lemma and deduce that there exists a sober space which is not strictly sober. Let $X$ be a set and let $\mathcal{P}(X)$ be its powerset, let $F\subseteq\mathcal{P}(X)$ be a filter such that it is not contained in any ultrafilter. Now, consider the Stone dual $X^{S}$ of $\mathcal{P}(X)$, let $\varphi$ be its topologizing map. Note that this is an isomorphism of Boolean algebras, as $\mathcal{P}(X)$ is atomic. All elements of the form $\varphi(Y)$ for $Y\subseteq X$ are clopens of the space $X^{S}$, hence compact. We then have that the filter of opens ${\uparrow}\varphi[F]$ is then Scott-open. By assumption, $\bigcap\varphi[F]=\emptyset$, and so if ${\uparrow}\varphi[F]$ is a neighborhood filter of some compact open this must be $\emptyset$, but this is not the case as the neighborhood filter of $\emptyset$ contains $\emptyset$, and $\emptyset\notin\varphi[F]$ by injectivity of $\varphi$. Thus, the space $X^{S}$ is sober, as it is a Stone space, but it is not strictly sober.
Lemma 5.6.
The Prime Ideal Theorem is equivalent to the statement that every Scott-open filter is an intersection of completely prime filters.
Proof.
We need to show that every Scott-open filter being an intersection of completely prime filters is equivalent to SPET. Suppose that SPET holds, and that $L$ is a frame and $F\subseteq L$ a Scott-open filter. Suppose, towards contradiction, that there is some $a\notin F$ such that $a\in P$ whenever $P$ is a completely prime filter with $F\subseteq P$. By SPET, there is a prime element $p\in L$ with $a\leq p$ and $p\notin F$. The completely prime filter $L{\setminus}{\downarrow}p$ contains $F$ but not $a$, and this is a contradiction. Conversely, suppose that every Scott-open filter is in $\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$. Let $F\subseteq L$ be a Scott-open filter, and suppose that $a\notin F$. There has to be a prime $p\in L$ such that $F\subseteq L{\setminus}{\downarrow}p$ and such that $a\notin L{\setminus}{\downarrow}p$. ∎
Proposition 5.7.
The following are equivalent.
  1. 1.
    The Prime Ideal Theorem holds.
  2. 2.
    We have $\mathsf{Filt}_{\mathcal{SO}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{CP}}(L))$ for every frame $L$.
  3. 3.
    $\mathcal{CP}$-compact Raney extensions are $\mathcal{SO}$-compact.
  4. 4.
    Sober spaces are strictly sober.
  5. 5.
    For a sober space $X$, the canonical extension of its frame of opens is $(\Omega(X),\mathcal{U}(X))$.
Proof.
That (1) and (2) are equivalent follows from Lemma 5.6. If (2) holds, (3) follows by the characterization in Proposition 3.2. Suppose, now, that (3) holds. For a sober space $X$, the Raney extension $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{CP}$-compact, by Proposition 3.11. Therefore, $(\Omega(X),\mathcal{U}(X))$ is $\mathcal{SO}$-compact, by hypothesis, hence strictly sober, by the characterization in Lemma 5.4. Items (4) and (5) are equivalent by Proposition 5.5. Finally, (4) implies (1) by Example 5.1. ∎
Finally, we give another proof, based on collections of filters, of the result in [bezhanishvili22] that the canonical extension of a Boolean algebra $B$ is the Booleanization of $\mathcal{U}(\mathsf{Idl}(B))$.
Proposition 5.8.
([jakl20], Proposition 8.1) For a coherent frame $L$, its canonical extension is the canonical extension of the distributive lattice $K(L)$ of its compact elements.
Lemma 5.9.
For a frame $L$, if $k\in L$ is a complemented element, then in the frame $\mathsf{Filt}(L)$ the filters ${\uparrow}k$ and ${\uparrow}\neg k$ are mutual complements.
Proof.
We have that $\neg{\uparrow}k=\{a\in L:a\vee k=1\}$, and this equals ${\uparrow}\neg k$. By definition of complement, also ${\uparrow}k\cap{\uparrow}\neg k=\{1\}$ and ${\uparrow}k\vee{\uparrow}\neg k=L$. ∎
Lemma 5.10.
For a compact, zero-dimensional frame $L$, Scott-open filters are exactly the joins of filters of the form ${\uparrow}k$, where $k\in L$ is a complemented element.
Proof.
Let $L$ be a compact, zero-dimensional frame, and let $F$ be a Scott-open filter. Let $f\in F$. We have $F=\bigvee_{i}k_{i}$, where $\{k_{i}:i\in I\}$ is the family of complemented elements below it. Observe that this is directed, and so there must be $j\in I$ with $k_{j}\in F$. For the converse, if $K$ is any family of complemented elements, suppose that there is a directed family $D\subseteq L$ such that $k\leq\bigvee D$ for some $k\in D$. As $L$ is compact, $k$ is compact, too, and so $k\leq d$ for some $d\in D$. Then, $\bigvee_{k\in K}{\uparrow}k$ is Scott-open. ∎
Lemma 5.11.
For a compact, zero-dimensional frame $L$, we have $\mathsf{Filt}_{\mathcal{R}}(L)=\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$.
Proof.
Because $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$, and this is the smallest sublocale containing $\{1\}$, for the inclusion $\mathsf{Filt}_{\mathcal{R}}(L)\subseteq\mathcal{I}(\mathsf{Filt}_{\mathcal{SO}}(L))$ it suffices to show that $\{1\}$ is Scott-open, but this follows immediately from compactness of $L$. For the other direction, it suffices to show that for a Scott-open filter $F$ we have $\neg\neg F\subseteq F$. Let $F$ be a Scott-open filter. By Lemma 5.10, this is $\bigvee_{i}{\uparrow}k_{i}$ for some collection $k_{i}\in L$ of complemented elements, which we can assume to be closed under finite meets without loss of generality. We have $\neg\neg\bigvee_{i}{\uparrow}k_{i}=\neg\bigcap_{i}\neg{\uparrow}k_{i}=\neg\bigcap_{i}{\uparrow}\neg k_{i}$, where we have used Lemma 5.9 for the last equality. Note also that $\neg\bigcap_{i}{\uparrow}\neg k_{i}=\neg{\uparrow}\bigvee_{i}\neg k_{i}$. Now, if $x\in\neg\neg F$ this means that $\bigvee_{i}\neg k_{i}\vee x=1$, and by compactness this means that $\neg k_{j}\vee x=1$ for some $j\in I$. Therefore $k_{j}\leq x$, and so $x\in F$. ∎
Proposition 5.12.
Let $L$ be a compact, zero-dimensional frame. Its canonical extension is
$(L,\mathsf{Filt}_{\mathcal{R}}(L)^{op}).$
This is also the canonical extension of the Boolean algebra $K(L)$.
Proof.
The first part of the claim follows from Lemma 5.11. The second part of the claim follows from Proposition 5.8, and the fact that $\mathsf{Filt}_{\mathcal{R}}(L)$ is the Booleanization of $\mathsf{Filt}(L)$. This is also the Booleanization of $\mathcal{U}(L)$, because for each $x\in L$ we have $\neg{\uparrow}x=\{y\in L:x\vee y=1\}$, and so the regular elements of $\mathcal{U}(L)$ are precisely the intersections of upsets of this form, that this, the regular filters of $L$. ∎
[…]